### v4

# May the Feedback Be with You! Breaking the Seal of Feedback-Driven Deep Learning Framework Fuzzing via Large Language Models

Abstract
Deep Learning (DL) frameworks have served as fundamental components in DL systems over the last decade. However, bugs in DL frameworks could lead to catastrophic consequences in critical scenarios. A simple yet effective way to find bugs in DL frameworks is fuzz testing (Fuzzing). Existing approaches focus on test generation, leaving execution results with high semantic value (e.g., coverage information, bug reports, and exception logs) in the wild, which can serve as multiple types of feedback.
To fill this gap, we propose FUEL to effectively utilize the feedback information, which comprises two Large Language Models (LLMs): analysis LLM and generation LLM. Specifically, analysis LLM infers analysis summaries from feedback information, while the generation LLM creates tests guided by these summaries. Furthermore, based on multiple feedback guidance, we design two additional components: (i) a feedback-aware simulated annealing algorithm to select operators for test generation, enriching test diversity. (ii) a program self-repair strategy to automatically repair invalid tests, enhancing test validity. We evaluate FUEL on the two most popular DL frameworks, and experiment results show that FUEL can improve line code coverage of PyTorch and TensorFlow by 4.48% and 9.14% over four state-of-the-art baselines. By the time of submission, FUEL has detected 104 previously unknown bugs for PyTorch and TensorFlow, with 93 confirmed as new bugs, 53 already fixed. 14 vulnerabilities have been assigned CVE IDs, among which 7 are rated as high-severity with a CVSS score of "7.5 HIGH".
1 Introduction
Deep Learning (DL) techniques have achieved success in various domains over the last decade, including healthcare [healthcare], chatbots [chatgpt], autonomous driving [huval2015autonomousdriving], and finance security [finance]. In these domains, popular DL frameworks like PyTorch [PyTorch] and TensorFlow [Tensorflow] help developers build, train, compile, and deploy DL models. Despite their widespread use, these frameworks remain susceptible to bugs, which can pose serious consequences, particularly in safety-critical applications [uberkill]. Hence, ensuring the quality of DL frameworks is essential, and various automated testing techniques [wang2020lemon, pham2019cradle, mu2024devmut, liu2023nnsmith] are proposed to detect bugs in DL frameworks. Among them, fuzzing [SuttonFuzzingBook, herrera2021seed, gong2025snowplow] is a prominent technique that systematically generates numerous inputs to uncover potential bugs. Traditional fuzzers [wang2020lemon, liu2023nnsmith] typically require manual design for mutation rules or operator specifications, which demands extensive domain knowledge and great human effort.
To alleviate the above issue, recent research has introduced Large Language Models (LLMs) [chatgpt, openai2023gpt4] into fuzzing, motivated by their success in other fields [SEsur]. One of the first classical studies is TitanFuzz [titanfuzz], whose key insight is that LLMs can implicitly learn both language syntax and intricate DL API constraints for DL program generation. Inspired by TitanFuzz, some follow-up studies [fuzzgpt, promptfuzz, whitefox, kernelgpt] are proposed to explore the potential of LLMs in fuzzing. The core idea behind these studies is to leverage the strong generative abilities of LLMs to generate or mutate inputs directly. Under such a trend, the “magnificent edifice” of fuzzing has already been built, but one dark cloud still hangs in the air: Have we fully exploited the potential of feedback information during the fuzzing loop, especially the semantically rich ones? Before answering this question, we first analyze the limitations of existing studies:
Figure 1: Motivation example of illustrating Limitation #1&2
Limitation #1: Overlooking multiple types of feedback. Previous approaches primarily focus on leveraging the powerful generation capabilities of LLMs to generate random inputs or edge cases. However, they tend to overlook the rich information derived from multiple feedback types. For example, TitanFuzz [titanfuzz] and FuzzGPT [fuzzgpt] do not utilize any feedback in their fuzzing loop, merely employing differential testing [evans2007difftesting] to identify potential bugs. Regrettably, the absence of feedback guidance leads to decreased code coverage and leaves more bugs undetected. In contrast, PromptFuzz [promptfuzz] employs a coverage-guided strategy to encourage LLMs to generate more diverse tests, alleviating the coverage plateaus problem. WhiteFox [whitefox] advances this idea by injecting compiler-optimization heuristics into the feedback loop. However, these techniques still only utilize a single and simple type of feedback during the whole fuzzing campaign and fail to consider multiple types of feedback simultaneously. As Figure 1 shows, each test execution yields three mutually exclusive feedback categories: coverage information, bug reports, and exception logs. However, how to fuse and exploit such complex feedback remains unexplored.
Limitation #2: Coarse-grained feedback information analysis. Each test execution in fuzzing yields various semantic information, which is usually a mixture of programming language and natural language. state-of-the-art (SOTA) feedback-driven fuzzers still distill feedback information into coarse signals (e.g., coverage increase), thereby discarding the fine-grained, context-rich semantic information. Even PromptFuzz [promptfuzz], which mitigates coverage plateaus, only focuses on whether new code is covered, ignoring semantics from concrete code snippets. This coarse-grained manner prevents the fuzzer from reasoning about the actual runtime state of the system under test (SUT). As Figure 1 shows, coverage information pinpoints specific code snippets, bug reports contain symptom descriptions, and exception logs provide detailed stack traces. Each feedback category can be effectively processed by LLMs, which transform raw feedback into actionable insights to guide subsequent test generation.
In addition to the utilization of feedback, existing techniques still have the following limitations in test generation.
Figure 2: Motivation example of illustrating Limitation #3
Limitation #3: Insufficient test diversity. TitanFuzz and FuzzGPT restrict themselves to API-level fuzzing (introduced in § 2.1), exhaustively enumerating isolated operators one by one. Although this paradigm achieves high API coverage, it overlooks bugs that manifest only when multiple operators interact within DL models. Model-level fuzzers such as WhiteFox [whitefox] synthesize tests that target compiler optimization patterns (e.g., operator fusion), yet this optimization-centric focus sacrifices test diversity, as many latent bugs lie outside the chosen optimization patterns. For instance, Figure 2 shows a test in which torch.frexp returns two tensors (torch.int32 and torch.float32) which are subsequently element-wise multiplied. Eager backend of PyTorch silently handles type promotion, whereas the compiler backend crashes with CppCompileError. TitanFuzz misses this bug due to their single-API testing paradigm. WhiteFox was also unable to detect this bug, as it can not generate such models that mismatch any optimization pattern. A natural and intuitive solution is to enable the LLM to consider more operator combinations by supplying external context information, thereby diversifying the generated tests.
Figure 3: Motivation example of illustrating Limitation #4
Limitation #4: Limited validity of generated tests. Even SOTA LLMs currently struggle to create valid tests for DL frameworks [deng2024can]. Existing techniques prompt LLMs with few-shot examples through in-context learning, but this paradigm poses limitations to LLMs’ understanding of the constraints on DL operators. Figure 3 shows an invalid test generated by WhiteFox that combines torch.mm and Add, imposing two constraints: (i) the first input of torch.mm must be column-aligned with the second; (ii) the inputs of Add must be either identically shaped or broadcastable. The generated tensors, however, have shapes $(2\times 10)$ and $(2\times 10)$, violating the above constraints and triggering a shape-mismatch error, which is a false positive. Our manual check reveals that 47.54% (290/610) of the tests generated by WhiteFox are invalid, similar to the above example. These failures not only waste fuzzing budgets but also lose opportunities to uncover bugs within operator interactions. Notably, we observe that most invalid tests are easy to fix. In this case, simply reshaping x1 from $(2\times 10)$ to $(2\times 2)$ generates a valid test that executes successfully and explores previously uncovered code paths. Equipped with the exception logs, LLMs are capable of automatically conducting such small-scale repairs, converting invalid executions into valid ones, thereby improving code coverage and bug detection capabilities.
Insight. LLMs have the potential to analyze multiple types of feedback information semantically and thereby generate diverse and valid tests. This motivates us to design two LLMs with distinct roles: (i) an analysis LLM distills complicated and redundant feedback information into concise summaries; (ii) a generation LLM synthesizes diverse and valid tests guided by analysis summaries during the fuzzing loop.
Our work. In this study, we propose FUEL, the first feedback-driven fuzzer that fully utilizes feedback information with two LLM-based agents, namely analysis LLM ($\mathcal{M_{A}}$) and generation LLM ($\mathcal{M_{G}}$). The key idea of our design is that LLMs could understand and analyze valuable feedback information. $\mathcal{M_{A}}$ first automatically infers the analysis summaries from the feedback. Then, analysis summaries are used to guide the $\mathcal{M_{G}}$ to generate tests that help to explore new code coverage or trigger new bugs (addressing Limitation #1 and Limitation #2). However, tests generated by LLMs usually contain common well-tested operators (e.g., Conv2d, ReLU), resulting in the lack of some uncommon poorly-tested operators (e.g., ConvTranspose, InstanceNorm). This behavior reduces test diversity. To solve this, we collect core operators from the DL framework API document as a retrieval set and design a feedback-aware simulated annealing algorithm to select operators from the retrieval set in each iteration. The retrieved operators, along with the analysis summary, are filled into the prompt of the $\mathcal{M_{G}}$ as context, allowing $\mathcal{M_{G}}$ to generate tests by using more operators (addressing Limitation #3). To improve test validity, FUEL introduces a program self-repair strategy to repair invalid tests. In simple terms, $\mathcal{M_{A}}$ in FUEL generates the repair strategy based on the exception log, which prompts the $\mathcal{M_{G}}$ to repair the test from the last iteration (addressing Limitation #4).
In summary, our work makes the following contributions:
  • Direction. To the best of our knowledge, we are the first to leverage LLMs to semantically analyze multiple types of feedback in fuzzing. Furthermore, beyond DL framework testing, this paradigm is also capable of conducting targeted fuzzing for other large and complex software systems by customizing more types of feedback (e.g., execution time and memory usage). Our promising results encourage further exploration in this direction.
  • Approach. We implement FUEL as a practical fuzzer for DL framework testing, which utilizes feedback information and repairs invalid tests. FUEL is available at https://github.com/NJU-iSE/FUEL.
  • Evaluation. Our experimental results show that the tests generated by FUEL achieve superior performance compared to SOTA baselines (NeuRI, TitanFuzz, NNSmith, and WhiteFox), improving line coverage on two DL frameworks (PyTorch and TensorFlow) by 4.48% and 9.14%, respectively. Additionally, ablation studies validate the contribution and effectiveness of each component in FUEL ’s design.
  • Real-World Contribution. So far, FUEL has detected 104 previously unknown bugs, with 93 already confirmed, 53 already fixed, and 14 assigned CVE IDs. 14 detected bugs were labeled as high priority, and one issue was labeled as utmost priority.
2 Background
2.1 DL Framework Testing
Previous studies on DL framework testing can be classified into API-level fuzzers [wei2022free, xie2022docter, titanfuzz] and model-level fuzzers [wang2020lemon, liu2023nnsmith, whitefox]. API-level fuzzers target testing each DL framework API by infilling its parameters. Notably, a DL framework API usually corresponds to a DL operator. For instance, torch.Conv2D, a PyTorch API (corresponding to 2-D convolution operator) can be filled with some parameters by API-level fuzzers, including kernel_size, stride, and padding. Then, a 4-D tensor will be created as input for torch.Conv2D. Finally, comparing outputs across different hardware (e.g., CPU and GPU) to detect potential bugs.
However, this paradigm is limited in its ability to detect bugs resulting from the intricate compositions of diverse operators in DL models. Therefore, model-level fuzzers are proposed to synthesize or mutate DL models that encapsulate a wide range of operators. A model-level test typically comprises two components: model definition and input tensors. As shown in Figure 3, the model definition is a Python class inheriting from the framework’s base Model class (e.g., torch.nn.Module). Meanwhile, the input tensors supply external tensor data to the model (e.g., torch.randn(4, 1)). Model-level fuzzers therefore strive to generate both the DL model (i.e., model definition) and input tensors that satisfy the model constraint, jointly forming a single model-level test. In this work, our designed FUEL is a model-level fuzzer, which aims to create diverse yet valid DL models.
Figure 4: Overview of FUEL
2.2 LLM-based Fuzzing
Until now, LLMs have achieved remarkable success in code intelligence tasks [hou2024llm4se]. Modern LLMs, typically based on Transformers, can be categorized into two types: Dense Transformer [transformer] and Mixture-of-Experts (MoE) architecture [jiang2024mixtral]. Dense Transformers (e.g., Llama-3 [dubey2024llama] and Qwen-2.5-Coder [hui2024qwen2.5-coder]) employ uniform parameter activation across all layers, enabling full contextual interaction through self-attention mechanisms, while MoE architecture models (e.g., Mixtral [jiang2024mixtral] and DeepSeek-V3 [deepseekai2024deepseekv3]) introduce sparse activation via expert routing algorithms, dynamically allocating computations to specialized subnetworks for improved scalability and efficiency. Recently, reasoning models (e.g., DeepSeek-R1 [deepseekai2025deepseekr1], GPT-o1 [openai-o1]) based on chain-of-thought (CoT) [wei2022cot] have attracted worldwide attention, showing impressive performance on math- and code-related benchmarks.
Impressive performance of LLMs motivates researchers to bring LLMs into software testing [wang2024software], especially fuzzing [llmfuzzsurvey]. In LLM-based fuzzing, LLM is usually employed as the test generator to synthesize programs. Along with instructions and the few-shot examples in prompts, it directly generates test code (similar to the code snippet in Figure 3). Specifically, these approaches primarily focus on using LLMs to generate or mutate inputs based on static prompts. However, they often overlook the execution results of generated tests, failing to utilize such valuable feedback to guide subsequent generation. In this context, FUEL augments the fuzzing loop with an analysis LLM to distill raw feedback information into actionable summaries, explicitly guiding the generation LLM to enrich test diversity and validity.
3 FUEL Design
Figure 4 depicts the end-to-end workflow of FUEL, integrating a $\mathcal{M_{A}}$ with a feedback-aware simulated annealing algorithm to utilize feedback information in a fully automated manner and subsequently combining analysis results to guide $\mathcal{M_{G}}$ to generate diverse and valid DL models. We show the details of each phase in the following subsections.
3.1 Prompting LLM for Test Generation
While generally employing a feedback-driven fuzzing approach, FUEL is designed to generate tests without feedback guidance in two key scenarios: (i) In the initial iteration of fuzzing: $\mathcal{M_{G}}$ performs purely stochastic generation based on predefined prompts, as currently no prior feedback exists for guidance. (ii) Upon model repair failure: FUEL incorporates a collaborative self-repair strategy (as introduced in § 3.3 and § 3.5) that engages both $\mathcal{M_{A}}$ and $\mathcal{M_{G}}$. When this self-repair fails in the current iteration, $\mathcal{M_{G}}$ temporarily disregards previous feedback and reverts to a non-guided generation in the next iteration cycle. The motivation behind this strategy is to prevent FUEL from being trapped in an infinite repair loop (i.e., attempt successive iterations persistently but fail to resolve preceding errors).
To accommodate these two scenarios, we design specialized default generation prompts that ensure continuous test generation even in scenarios where feedback is unavailable or repair stagnation occurs. The transition between feedback-driven and default generation is algorithmically controlled to optimize exploration-exploitation balance throughout the whole fuzzing campaign. Figure 5(a) shows a few-shot example in the default generation prompt. Firstly, it lists some operators selected by the heuristic search algorithm (discussed in § 3.4). Then this example goes with the instruction “Please generate a valid PyTorch model with selected operators”. Next, it provides a DL model consisting of two components: model definition and input tensors (introduced in § 2.1). Finally, with such few-shot examples, $\mathcal{M_{G}}$ generates tests with the prompt template shown in Figure 5(b). Placeholder [LIBRARY] in the instruction denotes the specific framework (PyTorch or TensorFlow), [SELECTED_OPS] is filled by results of heuristic search, and [PLACEHOLDER...] is the DL model to be generated by $\mathcal{M_{G}}$.
(a) A Few-Shot Example of PyTorch
(b) Default Prompt Template
Figure 5: Default prompt template of FUEL
3.2 Differential Testing and Oracles
Differential Testing. Differential testing has been proven to be an effective way to detect bugs in DL frameworks [liu2023nnsmith]. The DL model generated by LLM is executed on different backends to identify bugs. In this work, we focus on the compilation feature of DL frameworks and design a cross-check mode on two different backends as follows.
  • Eager backend refers to the DL framework that directly executes DL models without compilation. Following the development of frameworks over the past decade, eager backend has become stable and usually serves as a baseline for comparison with other backends.
  • Compiler backend has been the most important feature of DL frameworks since the version 2.x era, pushing the performance of DL frameworks to new heights. As an emerging technology, the compiler backend is still in rapid development, and has become an important backend to compile DL models (especially LLMs) [whitefox]. PyTorch Inductor and TensorFlow XLA are compiler backends for PyTorch and TensorFlow, respectively.
Oracles. We execute generated models on these two backends respectively and design two oracles to identify bugs.
  • Numerical inconsistency (silent incorrectness). The compiled model should be semantically equivalent to the native model (as the DL compiler only optimizes the performance of DL models). Otherwise, this implies that the compiler backend incorrectly optimizes the model, resulting in silent incorrectness. We use $|\text{input}_{i}-\text{other}_{i}|\leq atol+rtol\times|\text{other}_{i}|$ to compare two output tensors, where $\text{input}_{i}$ denotes each element in output tensor of eager backend and $\text{other}_{i}$ denotes each element in output tensor of compiler backend. We use the default tolerance (i.e., $atol=0.001$ and $rtol=0.001$).
  • Behavior inconsistency (program crashes). In addition to numerical inconsistency, aligning the behavior between the eager backend and the compiler backend is also very crucial. Both eager backend and compiler backend should behave the same on the same test inputs (i.e., both throw errors or pass tests). Otherwise, we consider this a potential bug (i.e., one throws errors, and the other one does not).
Execution and cross-checking on the two backends can be classified into three distinct states (Figure 1): (i) Cross-checking results violate the oracles, indicating a potential bug in DL frameworks, then we collect bug reports. (ii) If a model crashes on both backends, it is an invalid model, and we collect exception logs. (iii) Only when execution results of both backends do not violate oracles or crash, it is a bug-free and valid model, and code coverage is collected.
3.3 Prompting LLM for Analysis Summary
Figure 6: Analysis prompt templates of FUEL
As discussed above, test execution outcomes can be classified into three distinct states, and the post-execution phase following differential testing yields semantically rich feedback information. Our core insight lies in recognizing the potential of this semantic corpus (e.g., coverage data) as a valuable resource for LLMs. However, directly feeding the raw feedback into the generation LLM presents two critical challenges: (i) Lengthy and redundant feedback information. The inherent redundancy of code coverage contexts (typically spanning thousands of code blocks), coupled with verbose exception logs, imposes tremendous computational and understanding burdens on $\mathcal{M_{G}}$. (ii) Lack of strategic guidance. Unprocessed feedback fails to provide explicit and concise instructions for subsequent iterations, leaving $\mathcal{M_{G}}$ without guidance for generation.
To address the above challenges, we introduce a $\mathcal{M_{A}}$ to provide clear, explicit, and concise analysis summary for $\mathcal{M_{G}}$. $\mathcal{M_{A}}$ infers explanation, reasons, and next testing strategy from feedback step by step. Figure 6 shows the analysis prompt template of FUEL, which begins with a contextual description: “Shown here is a [LIBRARY] model along with input tensors from the last iteration.” The core components consist of [MODEL CODE], which represents a DL model (introduced in § 2.1). Next, $\mathcal{M_{A}}$ selects different templates based on different types of feedback, and the feedback information will be subsequently filled into the corresponding placeholders (i.e., [COVERAGE], [BUG], and [EXCEPTION]). Immediately following, the instruction makes the $\mathcal{M_{A}}$ analyze the feedback information semantically. Finally, $\mathcal{M_{A}}$ infers explanation, reasons, and next testing strategy in [PLACEHOLDER...]. This progressive step provides explicit guidance for test generation in subsequent iterations. More specific examples of analysis prompts will be discussed in § 5.4.
3.4 Heuristic Search for Operator Selection
While LLM-based approaches can benefit from sufficient analysis summaries, they tend to use some common and well-tested operators (e.g., Conv2d, ReLU, and Linear) for test generation [fuzzgpt], which significantly limits model diversity (discussed in Limitation #4). This limitation introduces the implementation of providing additional context information for $\mathcal{M_{G}}$, a logical and natural enhancement that prioritizes less prevalent operators in test generation.
Unfortunately, operator selection is a challenging NP-hard problem, stemming from the combinatorial explosion inherent in DL model construction: While DL frameworks typically offer a finite operator set (numbering in the thousands), the unbounded combinations of these operators create an intractably large retrieval space. Notably, we identify that feedback signals can serve as indicators of intrinsic value, enabling the iterative optimization of operator selection through value-guided search. The idea behind employing heuristic search lies in reformulating the operator selection task as a combinatorial optimization problem within a constrained space: At each iteration, the algorithm selects an optimal subset of operators from the pre-collected API documents that maximizes a predefined value function. This value function, dynamically determined by accumulated feedback signals, quantifies the strategic importance of operators for subsequent iterative generation.
Inspired by the above idea, we adopt simulated annealing [van1987simulated] for operator selection in FUEL, a balance between overhead and performance. High-performance search algorithms, such as Monte Carlo Tree Search, incur prohibitive temporal overhead during incremental tree construction in iterative fuzzing cycles. Conversely, low-overhead search algorithms (e.g., random sampling) fail to identify high-value operators, compromising search performance. Simulated annealing emerges as an offline planning algorithm that efficiently approximates global optima (regulated through cooling rate $\gamma$) within constrained time budgets. For operator valuation, we design three metrics of operators: (i) Number of uses; (ii) Exception-triggering count; (iii) Number of new triggered line coverage. The valuation function follows an inverse relationship with the first two factors (i.e., frequency and exceptions), reflecting our emphasis on exploring underused operators. Conversely, it maintains direct proportionality to new code coverage, prioritizing operators that can trigger new code coverage. This strategy effectively prevents the fuzzer from being trapped in coverage plateaus caused by over-testing common operators. We define the final valuation function as follows: $V_{\alpha,\beta}(x,y,z)=\frac{\alpha}{\alpha+x}+\frac{\alpha}{e^{y}}+\beta-e^{-\frac{z}{100}}$, where $x$, $y$, and $z$ denote for number of use, exception-triggering count, and number of new triggered line coverage, respectively. The hyper-parameters $\alpha$ and $\beta$ are both initialized at $0.5$. This configuration ensures algorithmic stability before the fuzzing campaign, where $x$, $y$, and $z$ remain zero-initialized. Under this initialization setting, the value for all operators reduces to: $V_{\alpha,\beta}(0,0,0)=\alpha+\beta$. Substituting the predefined hyperparameters yields a uniform initial value $V_{0.5,0.5}(0,0,0)=1.0$, ensuring fair operator prioritization at the beginning of fuzzing.
Algorithm 1 presents the details of our Feedback-Aware Simulated Annealing (FASA). The algorithm begins by verifying whether an exception was triggered in the previous iteration (Line 1). If true, FASA returns the previously selected operator sequence to provide a repair opportunity for FUEL (Lines 1-1). Subsequently, operator values are updated through the integration of three distinct feedback signals (Lines 1-1). The optimization process employs a simulated annealing framework that returns the selected operators (Lines 1-1), initialized with standard temperature parameters and cooling schedules (Lines 1-1). During each annealing cycle (Line 1), $N_{s}$ iterations are performed to evaluate candidate operator sequences $S_{new}=[op_{1},op_{2},...,op_{k}]$ (Lines 1-1), where the composite fitness value function $F$ is calculated as the average of individual operator values ($F=\frac{1}{k}\sum_{i=1}^{k}V_{\alpha,\beta}(x_{i},y_{i},z_{i})$). While sequences with $\Delta F\geq 0$ are typically accepted, the Metropolis criterion probabilistically permits inferior solutions (with probability $e^{\Delta F/T}$) to escape local optima (Lines 1-1). The temperature parameter undergoes reduction according to a predefined cooling schedule until reaching termination thresholds (Lines 1-1). Finally, FASA returns a selected operator sequence in the current fuzzing iteration (Line 1).
[…]
 Output : Selected operator sequence: selectedOps
 if exceptionFlag ==1==1 then
    selectedOps $\leftarrow$ lastOps
    return selectedOps
   end if
 if isNotFirstIteration then
    for $\textnormal{{Op}}\in\textnormal{{lastOps}}$ do
       $\textnormal{{Op}}_{used\_times}\leftarrow\textnormal{{Op}}_{used\_times}+1$
       $\textnormal{{Op}}_{exp\_count}\leftarrow\textnormal{{Op}}_{exp\_count}+$ exceptionFlag
[…]
      end for
   end if
 selectedOps $\leftarrow$ SimulatedAnnealing(opSet)
 return selectedOps
Function SimulatedAnnealing(opSet):
 $T,N_{s},\gamma$ $\leftarrow$ $100,10,0.99$
 $k$ $\leftarrow$ RandomNumber($1,3$)
 $S_{cur}$ $\leftarrow$ RandomSelect($\textnormal{{opSet}},k$)
 while T $\geq$ $T_{min}$ do
    $n_{s}$ $\leftarrow$ 0
    while $n_{s}$ < $N_{s}$ do
       $S_{new}$ $\leftarrow$ RandomSelect($\textnormal{{opSet}},k$)
[…]
          $S_{cur}$ $\leftarrow$ $S_{new}$
         end if
       $n_{s}\leftarrow n_{s}+1$
      end while
    $T\leftarrow T\cdot\gamma$
   end while
 return $S_{cur}$
Algorithm 1 Feedback-Aware Simulated Annealing
3.5 Feedback-guided Prompting for Test Generation
Figure 7: Feedback-guided prompt template of FUEL
Feedback guidance consists of the analysis summaries generated by $\mathcal{M_{A}}$ and selected operators searched by FASA. FUEL employs three testing strategies (i.e., explore new coverage, trigger similar bugs, and repair invalid tests). As illustrated in Figure 7, feedback-guided prompts first specify generation instructions and the DL model generated in the last iteration, followed by the interaction of [ANALYSIS SUMMARY] containing next testing strategy. Meanwhile, [SELECTED_OPS] proposes selected operators. The [PLACEHOLDER...] serves as the target position for the new DL model generated by $\mathcal{M_{G}}$. Notably, feedback-guided prompts operate mutually exclusively with the default prompt described in § 3.1. Activation conditions of feedback-guided prompts require the simultaneous satisfaction of two criteria: (i) the current iteration is not the first iteration. (ii) No consecutive invalid executions have been detected in recent cycles. This conditional substitution mechanism ensures adaptive utilization of historical feedback while maintaining generation stability, with the feedback-guided prompt dynamically replacing the default template. This design enables FUEL to effectively balance exploratory test generation with feedback-driven refinement through its context-sensitive prompt switching strategy.
4 Implementation
Operator collection from API document. As discussed in § 2.1, an operator set is needed for FASA. In this work, we collect core operators from two DL frameworks. Specifically, we implement an HTML crawler to collect operators from PyTorch with API prefixes torch, torch.nn, and torch.nn.functional as they are core operators [ptop] in PyTorch 2.x. For TensorFlow, we collect operators that XLA supports from its document [tfxlaops] as we focus on the compilation feature of the DL framework in this work. Although we currently only consider the most important subset of all operators, we can easily modify the operator set to suit specific test scenarios (e.g., considering gradient-related operators to test the gradient feature of DL frameworks).
Generation and analysis LLMs. While closed-source models (e.g., GPT-4o [gpt-4o]) demonstrate amazing performance, their costs are substantial through the API service. In contrast, open-source models enable local deployment but typically exhibit comparatively inferior performance. In our approach, we use DeepSeek-V3 [deepseekai2024deepseekv3] as the $\mathcal{M_{A}}$ for feedback analysis, leveraging its demonstrated superiority in most benchmarks related to text summary. Meanwhile, we select Qwen-2.5-Coder-32B [hui2024qwen2.5-coder] as $\mathcal{M_{G}}$ for test generation due to its leading performance on code generation benchmarks [evalplus]. The temperature of the Qwen model is set to one that aligns with the mechanism of fuzzing (i.e., generating random and different test cases each time). This model selection strategy carefully balances the trade-off between performance and costs. Specifically: (i) For local deployment, we use vLLM [vllm], an efficient LLM inference engine with optimized memory management, to host open-source local models; (ii) For API services, we adopt a Python package openai [openai-python] that maintains compatibility with mainstream model API services. Notably, our artifact supports flexible integration of diverse LLMs through different invocation ways (i.e., local deployment and API services), allowing dynamic selection between generation and analysis models based on specific requirements and budgets.
In-Context Learning Prompting. Prompt engineering plays a pivotal role in LLM performance [sahoo2024promptsurvey]. Since FUEL integrates two distinct LLM agents ($\mathcal{M_{A}}$ and $\mathcal{M_{G}}$), we design specialized prompts for each, employing lightweight few-shot In-Context Learning (ICL) [brown2020gpt3]. To balance manual effort with system extensibility, we curate a concise set of few-shot examples that comprehensively cover the spectrum of feedback categories defined in our design: (i) Coverage Information, distinguishing whether triggering new coverage; (ii) Bug Reports, encompassing both single-backend crashes and numerical inconsistencies; (iii) Exception Logs, representing invalid executions where crashes occur across both backends. To construct high-quality examples, we collected representative DL programs from framework documentation and historical GitHub issues. These programs were executed to capture real-world <program, feedback> pairs. We then manually crafted the corresponding [ANALYSIS SUMMARY] to form ground-truth triplets. These triplets serve as few-shot examples to teach $\mathcal{M_{A}}$ how to distill feedback, and subsequently guide $\mathcal{M_{G}}$ in interpreting these summaries for test generation. Although we opt for this lightweight prompting strategy over complex paradigms (e.g., Chain-of-Thought [wei2022cot]) for efficiency, it demonstrates sufficient effectiveness in our evaluation. Notably, such implementation significantly reduces the manual burden compared to traditional fuzzing techniques [liu2023nnsmith, wang2020lemon].
5 Evaluation
5.1 Research Questions
We design the following research questions (RQs) in our evaluation:
  • RQ1: Can FUEL outperform the state-of-the-art (SOTA) baselines on coverage and bug detection?
  • RQ2: Can FUEL detect real-world new bugs?
  • RQ3: How does each component contribute to FUEL?
5.2 Experimental Setup
Systems under test. In this work, we focus on two most popular DL frameworks (PyTorch [PyTorch] and TensorFlow [Tensorflow]). The statistics of two DL frameworks are shown in Table 1. In the following tables, we use PT and TF to refer to PyTorch and TensorFlow, respectively.
Table 1: Statistics of the SUTs
SUT Lines of Code Test Language Tested Version
PT 322,839 Python Nightly20250214
TF 285,406 Python Nightly20250228
Fuzzing configuration and environment. We run FUEL on 1,000 iterations to generate tests, as a trade-off between high costs and the number of sufficient generated tests. Our experiments were conducted on Ubuntu 20.04 LTS with 4 NVIDIA V100 GPUs.
Baselines. We select SOTA DL framework testing baselines from two categories (LLM-based fuzzers [titanfuzz, whitefox] and traditional fuzzers [liu2023nnsmith, liu2023neuri]). To ensure a fair comparison, we use the same oracle and differential testing (as detailed in § 3.2) for baselines. We introduce the characteristics of these baselines as follows.
  • TitanFuzz [titanfuzz] is the first technique to directly leverage LLMs for fuzzing DL libraries, capitalizing on their implicit knowledge of language syntax and API constraints. It employs a generative LLM to produce seed programs and an infilling LLM to populate valid API parameters. Finally, it executes generated APIs and compares outputs across different hardware to detect bugs.
  • WhiteFox [whitefox] is an LLM-based white-box fuzzer that leverages compiler optimization source code to guide test generation. It uses an analysis LLM to extract optimization patterns as requirements, which are then used by a generation LLM to synthesize targeted test cases.
  • NNSmith [liu2023nnsmith] targets DL compilers by utilizing an SMT solver with manually defined operator specifications to generate diverse and valid computation graphs. Then it further employs a gradient-based search to find inputs that prevent floating-point exceptions, ensuring valid comparisons during differential testing.
  • NeuRI [liu2023neuri] tackles the challenge of defining complex operator constraints by employing inductive program synthesis. It infers these constraints automatically from collected valid and invalid API traces and subsequently utilizes a hybrid model generation strategy (combining symbolic and concrete execution) to synthesize diverse and valid DL models capable of supporting hundreds of operator types.
FUEL variants. We design five variants for ablation studies.
[…]
Metrics. We consider two types of metrics. (i) Code coverage: We use a Python library (namely Coverage.py [coverage-py]) to collect line code coverage. (ii) Number of detected bugs: Notice that we also record the number of detected bugs. We report our detected bugs to the corresponding repository issues on GitHub. We record the issue as “Confirmed” only when the developer can reproduce it and as “Fixed” when a pull request (PR) linked to the issue is merged into the main branch.
5.3 RQ1: Comparison with SOTA baselines
5.3.1 Default fuzzing configuration
Figure 8 demonstrates the performance of FUEL against four SOTA baselines, on both PyTorch and TensorFlow. Over 1,000 iterations, FUEL achieves superior code coverage compared to all baselines, outperforming the second-best baseline NeuRI by 4.48% and 9.14% on PyTorch and TensorFlow, respectively. Among the baselines, NeuRI demonstrates the most competitive performance. Although NeuRI is also a model-level fuzzer capable of synthesizing valid models via inductive rule inference, it is fundamentally limited by its reliance on random operator combinations. This stochastic nature prevents it from systematically exploring the rich and deep interaction space between operators. In contrast, FUEL utilizes feedback-driven guidance and heuristic search to strategically combine operators, thereby unlocking harder-to-reach code paths that random combinations miss. Other baselines show distinct limitations. For instance, WhiteFox exhibits relatively lower coverage, potentially attributed to its optimization-triggering pattern specialization. While effective for testing specific compiler optimizations, this feature inadvertently restricts model diversity through excessive template matching. These experimental results support our hypothesis that utilizing coverage information to guide test generation is more effective than LLM-based fuzzers and traditional fuzzers, achieving greater coverage across 1,000 iterations.
(a) PyTorch
(b) TensorFlow
Figure 8: Code coverage comparison with SOTA baselines.
5.3.2 24-hour fuzzing configuration
Following prior studies [xia2024fuzz4all, whitefox], we evaluate FUEL over a 24-hour fuzzing campaign. Table 2 reveals that FUEL generates the fewest tests (approximately 8.80% of TitanFuzz’s tests) on PyTorch. This phenomenon is attributed to the computationally intensive nature of the multi-agent architecture within our fuzzing loop. Despite the significant overhead, FUEL shows competitive code coverage compared to TitanFuzz and, more importantly, superior bug detection capabilities. Notably, TitanFuzz failed to detect any new bugs on the latest versions of PyTorch and TensorFlow. Similarly, NeuRI generates the most tests (588,610 on PyTorch and 649,924 on TensorFlow) by leveraging efficient inductive rule inference and its own engineering optimization. However, its reliance on random operator combinations tends to produce numerous simple or repetitive tests. Consequently, NeuRI detects fewer bugs and achieves lower code coverage than FUEL. These results indicate that in DL framework testing, simple and extensive test generation (i.e., generating a single API or random operator combination) provides small rewards, especially in bug detection, emphasizing the need for strategic generation of architecturally diverse models rather than quantity-driven approaches. In terms of model validity and diversity, the near-perfect validity of TitanFuzz stems from its API-level fuzzing paradigm that inherently bypasses the complex constraint verification required for operator interactions. Among model-level testing approaches, NNSmith achieves 100% validity rate by design as it utilizes SMT solvers to construct computation graphs. However, it relies on a small set of manually defined operators that are thoroughly tested, which restricts its ability to explore more unusual operators. FUEL demonstrates significant validity improvements over WhiteFox, achieving 24.82% and 41.63% higher validity rates in PyTorch and TensorFlow, respectively. This performance gap highlights FUEL’s enhanced capability in repairing invalid models through its self-repair strategy. The 24-hour experimental results position FUEL as the balanced solution that addresses the trade-off between model validity and diversity in DL framework fuzzing.
Table 2: Comparison with baselines over 24-hour fuzzing
. SUT Technique # Bugs Coverage # Valid Tests (%) # Tests PT FUEL 8 97,195 2,808 (78.28%) 3,587 TitanFuzz 0 98,487 36,734 (90.21%) 40,721 WhiteFox 1 88,390 3,944 (51.55%) 7,651 NNSmith 1 87,922 103,248 (100.00%) 103,248 NeuRI 2 92,274 588,610 (93.09%) 632,310 TF FUEL 2 94,114 1,930 (64.03%) 3,014 TitanFuzz 0 99,778 33,896 (93.84%) 36,121 WhiteFox 0 74,785 1,353 (22.40%) 6,041 NNSmith 0 75,777 47,321 (100.00%) 47,321 NeuRI 0 76,493 649,924 (89.72%) 724,392
5.4 RQ2: Bug Study
Table 3 presents the bug detection results of FUEL. We first analyze some representative detected bugs and then analyze the advantage of FUEL to trigger similar new bugs. Lastly, we explain why some detected bugs won’t be fixed by DL framework developers.
Table 3: Summary of our detected bugs
SUT Total Confirmed (Fixed) Pending Rejected
PT 98 87 (53) 6 5
TF 6 6 (0) 0 0
Total 104 93 (53) 6 5
5.4.1 Analysis of bug examples
As discussed in § 3.2, we use two oracles to identify bugs on two backends, including numerical inconsistency and behavior inconsistency. We introduce two representative bugs from PyTorch and TensorFlow as follows.
(a) A PyTorch numerical inconsistency bug detected by FUEL
(b) A TensorFlow behavior inconsistency bug detected by FUEL
Figure 9: Representative bugs detected by FUEL
Figure 9(a) shows a numerical inconsistency bug of PyTorch which implements an incorrect graph-level optimization. The two tensors returned by torch.chunk are subtracted, resulting in different outputs on the two backends. The root cause of this bug is that PyTorch Inductor attempts to fuse torch.nn.ReflectionPad3d, view, and torch.chunk in the computational graph (this optimization is also called operator fusion [alwani2016fused]). However, in the final step of indexing (i.e., x[0]- x[1]), PyTorch Inductor incorrectly assumed that the value of the tensor must be positive (while it might be zero due to the modulo operation), causing the compiler backend to output incorrect results. This bug was labeled as utmost priority by PyTorch community and assigned CVE-2025-46148, which was fixed by a PR immediately.
Figure 9(b) demonstrates a behavior inconsistency bug of TensorFlow, which is a security vulnerability detected by FUEL. The buggy model consists of tf.reshape and tf.keras.layers.Conv2D. When setting padding='valid' for Conv2D, the eager backend executes the model successfully, while the compiler backend throws negative dimension size error. This vulnerability can be leveraged by requiring the Conv2D parameter in padding='valid' format, resulting in DoS attacks, which was assigned CVE-2025-55559. CVSS-3.x scored the severity of this vulnerability as "7.5 HIGH".
(a) Trigger a new similar bug through operator overloading
(b) Trigger a new similar bug through parameter mutation
(c) Trigger a new similar bug through inverse operators
Figure 10: The process of FUEL triggering new similar bugs with the last triggered bugs
5.4.2 Analysis of detecting similar new bugs
In our approach, one of the feedback goals is to detect new bugs with the last triggered bug. The insight behind this goal is that bug-inducing inputs are usually clustered together [dickinson2001pursuingfailure]. $\mathcal{M_{A}}$ can analyze current bugs and guide $\mathcal{M_{G}}$ to generate similar bug-inducing inputs in the next iteration. FUEL can generate similar bug-inducing inputs from three perspectives, including operator overloading, parameter mutation, and inverse operators.
Operator overloading. Figure 10(a) illustrates the process of generating bug-inducing inputs. First, the last test triggers a bug with the FractionalMaxPool2d operator, an optimization problem with inconsistent results. Then, the analysis of the current test results is given by $\mathcal{M_{A}}$. It highlights that this could be a common issue with the FractionalMaxPool and suggests that test generation uses the FractionalMaxPool3d operator in the next iteration. The $\mathcal{M_{G}}$ generates the test, which is executed on different backends based on the analysis summaries. The inconsistent results after execution indicate that FractionalMaxPool3d was also incorrectly optimized by PyTorch Inductor. We reported these two bugs to PyTorch and a core developer commented on them as follows:
“Thank you for the issues. Silent incorrectness is always a high priority!” — A PyTorch core developer
Parameter mutation. In addition to detecting bugs of operator overloading, FUEL can also detect bugs for different parameters (i.e., data types and properties) of the same operator. Figure 10(b) shows two bugs found by FUEL. Feedback from the last iteration is a model consisting of torch.argsoft and F.avg_pool2d, triggering numerical inconsistency. $\mathcal{M_{A}}$ infers that return value of torch.argsoft is torch.int64 which results to all-zero output of F.avg_pool2d on inductor. Hence, $\mathcal{M_{A}}$ suggests using other unusual data types (e.g., torch.uint64) to test F.avg_pool2d in the next iteration. $\mathcal{M_{G}}$ generates the test based on the suggested strategy and subsequently triggers a behavior inconsistency bug. PyTorch developers confirmed this bug, and we sent a PR to fix it, which was quickly approved and merged into main branch.
Inverse operators. Different from the above two perspectives, which trigger similar bugs on the same operator, FUEL can also target different operators. As Figure 10(c) shows, in the last execution, bitwise_right_shift outputs incorrectly on the compiler backend. $\mathcal{M_{A}}$ infers that an inverse operator (i.e., bitwise_left_shift) may have the same bug and suggests testing this operator in the next iteration. $\mathcal{M_{G}}$ subsequently generates the DL model consisting of bitwise_left_shift which triggers the silent incorrectness similar to bitwise_right_shift. This bug was labeled as high priority and assigned CVE-2025-46152.
5.4.3 Rejected Bugs
Among the 104 bugs we have reported, five of them won’t be fixed by developers. Developers dismiss them because the numerical inconsistency is too small. They suggested increasing the tolerance threshold of accuracy. Subsequently, we did it in our next bug reports.
5.5 RQ3: Ablation Studies
Similar to the observations of prior studies (FuzzGPT [fuzzgpt], WhiteFox [whitefox]), we notice that the PyTorch community is more active than the TensorFlow community as PyTorch developers confirm and fix issues faster. Therefore, we run our ablation experiments on PyTorch with the default fuzzing configuration. Table 4 shows ablation comparison results.
Table 4: Comparison with ablation variants
[…]
5.5.1 Contribution of analysis LLM
[…]
5.5.2 Contribution of FASA
[…]
5.5.3 Contribution of different feedbacks
[…]
6 Discussion
6.1 Sensitivity Analysis of LLM Backends
To evaluate the sensitivity of FUEL on different LLM backends, we experiment with three combinations of $\mathcal{M_{A}}$ and $\mathcal{M_{G}}$. In addition to default implementation (§ 4), we consider two additional LLM pairs. (i) Qwen-2.5-Coder-32B [hui2024qwen2.5-coder] paired with Qwen3-30B [qwen3]: These two open-source LLMs were selected due to their compatibility with our local environment, and this setting allows us to examine how much performance degradation FUEL would experience in a zero-cost budget scenario. (ii) Claude-Sonnet-4 [claude-4] paired with GPT-5 [gpt-5]. These LLMs were chosen based on their SOTA performance on coding benchmark [swebench], and this setting allows us to investigate the performance gains when inspiring the maximum potential of FUEL.
[…]
Table 5: Performance of FUEL across different LLM backends
$\mathcal{M_{G}}$ $\mathcal{M_{A}}$ # Bugs Coverage \$Costs
Qwen-2.5-Coder-32B DeepSeek-V3 6 96,405 \$1.68
Qwen-2.5-Coder-32B Qwen3-30B 2 89,244 \$0.00
Claude-Sonnet-4 GPT-5 8 100,894 \$9.63
The above results indicate that FUEL is moderately sensitive to the choice of LLM backend. Lower-tier open-source LLMs tend to generate less diverse and valid tests and therefore detect fewer bugs. High-end frontier LLMs can better leverage feedback and thus improve the quality of both analysis and generation, although the improvements come at a substantially higher monetary cost. Such moderate sensitivity suggests that FUEL can maintain competitive performance even when deployed with mid-tier LLMs, while still retaining the ability to benefit from stronger LLMs when available. Notably, as future LLMs offer stronger reasoning, higher code generation, and more reliable feedback summarization ability, FUEL can directly inherit these advancements without requiring re-implementation. As a result, FUEL’s performance can be further enhanced with the progress of LLMs.
6.2 Limitations
FUEL incurs inherent computational costs due to the use of an LLM-based multi-agent architecture within the fuzzing loop, which results in notable challenges in both time and monetary overhead. (i) Time cost arises from the limited throughput of test generation in FUEL. As detailed in § 5.3.2, $\mathcal{M_{A}}$ must process a large amount of feedback in each iteration, which restricts the number of generated tests. Our current implementations demonstrate an average latency of 24.09 seconds per iteration. Fortunately, this limitation is expected to be alleviated by the rapid evolution of LLM inference optimization techniques [vllm]. (ii) monetary cost is introduced by the token-based billing model of commercial LLM API services, especially because feedback analysis consumes a substantial number of tokens. To reduce this expense, our implementation relies on cost-effective API providers such as the DeepSeek API [deepseek-api]. The complete experimental evaluation conducted using the DeepSeek API incurred a total cost of approximately \$40.89, corresponding to roughly \$0.39 per detected bug in our experiments.
7 Related Work
7.1 LLM-Based Fuzzing
TitanFuzz [titanfuzz] is one of the pioneering works in LLM-based fuzzing, which leverages two LLMs to generate and infill DL library APIs, respectively. Subsequently, Fuzz4All [xia2024fuzz4all] believes LLMs have been trained on the multilingual corpus to serve as a universal fuzzer for the different SUTs with different programming languages. Sedar [fu2024sedar] leverages LLMs to generate high-quality initial seeds by adapting test cases from other Database Management Systems (DBMSs), improving code coverage in target DBMS fuzzing. LLMIF [wang2024llmif] enhances IoT protocol fuzzing by using LLMs to extract protocol information and reason about device responses from protocol specifications. ChatAFL [meng2024large] utilizes LLMs to transform human-readable protocol specifications into machine-readable formats, enabling the fuzzer to generate and mutate protocol messages systematically. PromptFuzz [promptfuzz] employs coverage-guided prompt fuzzing techniques to generate fuzz drivers iteratively. More recently, KernelGPT [kernelgpt] automates the generation of Syzkaller specifications using LLMs, leveraging their extensive pre-training on kernel code and documentation to create valid syscall sequences. Unlike previous LLM-based fuzzers, FUEL is the first fuzzing technique to leverage an analysis LLM to analyze various semantic information in the feedback and achieve promising performance.
7.2 DL Framework Testing
In recent years, DL framework testing has become an active topic, and a large number of fuzzers have emerged, which can be classified into traditional fuzzers [pham2019cradle, wei2022free, xie2022docter, deng2022fuzzing, yang2023fuzzing, wang2020lemon, guo2020audee, gu2022muffin, liu2023nnsmith, liu2023neuri, mu2024devmut] and LLM-based fuzzers [titanfuzz, fuzzgpt, whitefox].
One of the pioneering works is CRADLE [pham2019cradle], which detects inconsistencies by running pre-built DL models across multiple low-level backends of Keras [Keras2020]. To enhance the diversity of models, LEMON [wang2020lemon] and AUDEE [guo2020audee] build upon CRADLE by applying pre-defined mutation rules to seed models and inputs, generating a broader range of test cases. NNSmith [liu2023nnsmith] models each DL operator with input constraints and shape inference, using SMT solving to generate valid models. Subsequently, NeuRI [liu2023neuri] is proposed to synthesize DL models via rule inference automatically. Unfortunately, the above traditional fuzzers need manual design for mutation rules, operator specifications, or synthesis rules, which need domain-specific expert knowledge. With the success of LLMs in software testing [lemieux2023codamosa, feng2024prompting], researchers have attempted to apply LLMs to fuzzing DL frameworks. TitanFuzz [titanfuzz] is the first work to leverage LLMs to generate API call sequences directly. Then FuzzGPT [fuzzgpt] is proposed to generate edge-case code via in-context learning and fine-tuning. WhiteFox [whitefox] proposes a white-box compiler fuzzing technique aided by LLMs, targeting the generation of specific inputs to trigger optimization features in compilers. However, all of them struggle to generate diverse and valid DL models, and FUEL is proposed to fill this gap by adopting a heuristic search for operator selection and a program self-repair strategy by mining the potential of feedback.
8 Conclusion
In this work, we propose FUEL, the first feedback-driven fuzzer for DL frameworks, which leverages LLMs to process feedback information in the fuzzing loop. FUEL employs an LLM-based multi-agent architecture comprising an analysis LLM and a generation LLM: the former infers concise analysis summaries from feedback information to guide the latter in generating tests that help to improve code coverage and bug detection. Moreover, based on multiple feedback guidance, FUEL adopts a heuristic search algorithm and program self-repair strategy to enhance model diversity and validity, respectively. Our extensive experiments validate the effectiveness of each component and demonstrate FUEL’s strong capability in detecting bugs and improving code coverage compared to SOTA baselines.
Acknowledgment
We thank Wei Cheng, Yanzhou Mu, and Zhiyuan Li for their insightful feedback on the manuscript.
Appendix A Failure Case Analysis
Figure 11: A failed self-repair of FUEL
As discussed in our ablation study, FUEL can repair most of the invalid models. However, not all the invalid models can be repaired. After our manual check, we find that failed repair attempts typically involve programs with excessive operators and complex constraints, where LLMs erroneously apply shape-conversion operators (e.g., View and Unsqueeze), zooming shape errors. Figure 11 presents a representative failure mode involving operator compatibility. The original model defines a Linar operator expecting an input dimension of 10, but receives a tensor x with shape (2, 50). Triggered by the dimension mismatch error, FUEL attempts to repair the invalid model by reshaping x into (10, 10) using additional .view(). While this satisfies the Linear operator’s constraint (producing an output shape (10, 20)), it implicitly destroys the batch dimension semantics. This “specious” fix causes a crash in the subsequent residual connection out + x, where the shapes (10, 20) and (2, 50) become irreconcilable. Finally, the patched model is still invalid, and both backends throw RuntimeErrors. A promising future direction is that LLM generates shape-agnostic DL programs (i.e., fuzz drivers) while leveraging constraint-solving techniques (e.g., Z3Prover [z3-solver]) to automatically resolve tensor shape and API parameters, thereby producing valid and diverse DL models through hybrid neuro-symbolic approaches.